OpenStudy (anonymous):

If f(x) = x^2 - √x, find f '(x).

OpenStudy (anonymous):

please help

myininaya (myininaya):

f(x)=x^2-x^(1/2) f'(x)=2x-1/2*x^(1/2-1)=2x-1/2*x^(-1/2)=2x-1/(2sqrt(x))